Air flows through a converging-diverging nozzle/diffuser. Assuming isentropic flow, air as an ideal gas, and constant specific heats determine the state at several locations in the system. Solve using equations rather than with the tables. Note: The specific heat ratio and gas constant for air are given as k=1.4 and R-0287 kJ/kg-K respectively --Given Values-- Inlet Temperature: TI (K) 349 Inlet pressure: Pl (kPa) 460 Inlet Velocity: V1 (m/s) 73 Area at nozzle inlet: Al (cmA2) 8.19 Throat area: A (cm 2) 2.88 Area at diffuser exit: A (cmA2) 6.33
Answer #1

Find the Mach number, 73 V1.4x 287 x 349 = 0.1949 (d) Find the theoretical throat area for sonic flow by using the following

- (14 24-70.199) PO = 1.0268 Po = 472.328kPa Find the pressure at the throat. P2 472.328 - 1.4387 P2 B2 = 328.301 kPa Find th